crude approximation of voice production is to consider the breathing passages and mouth to be a esonating tube closed at one end. What is the fundamental frequency if the tube is 0.240 m long, by taking air temperature to be 37.0°C? Number 367 Hz What would this frequency become if the persons breathing passage was filled with carbon dioxide instead of air? Assume the same temperature dependence for carbon dioxide as for air Number 1070 Hz
Answer #1

A)

Velocity of sound, v = 331 sqrt(310/273) = 352.71 m/s

Frequency, f = 352.71/(4 x 0.24) = 367 Hz

B)

Velocity, v = sqrt(1.28 x 8.314 x 310/0.044) = 274 m/s

Frequency, f = 274/(4 x 0.24) = 285.42 Hz

Comment in case any doubt please rate my answer ..